OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3). (explain help please)

OpenStudy (anonymous):

y = 2x + 6 y = −1/2x + 3 y = −1/2x + 6 y = 2x + 3

OpenStudy (anonymous):

so noone wants to help? thankss everyone

OpenStudy (shamim):

ur given equation of line is y=2x+2

OpenStudy (shamim):

i m now helping u

OpenStudy (shamim):

so the slope of ur given line is m=2

OpenStudy (shamim):

u know \[m _{1}m _{2}=-1\]

OpenStudy (shamim):

when 2 lines r perpendicular with each other

OpenStudy (shamim):

so the slope of ur required line will b 2m=-1

OpenStudy (shamim):

m=-1/2

OpenStudy (shamim):

r u following me

OpenStudy (anonymous):

y=2x+2 is the given equation here the slope = 2 now the line which will be perpendicular to this line will have the slope = -1/2 now we got the slope the point is (6,3) using point slope form we can get the equation y-y1 = m(x-x1) where m=-1/2 x1 = 6 , y1 = 3 y-3 = -1/2(x-3)

OpenStudy (shamim):

so the equation of ur required line will \[y=-\frac{ 1 }{ 2 }x+c\]

OpenStudy (anonymous):

sorry its y-3 = -1/2(x-6) => y= -1/2x +6 -----> required equation

OpenStudy (shamim):

or,\[3=-\frac{ 1 }{ 2 } \times 6 +c\]

OpenStudy (shamim):

because ur required line will go through the point 6,3

OpenStudy (shamim):

c=6

OpenStudy (shamim):

\[y=-\frac{ 1 }{ 2 }x+6\]

OpenStudy (shamim):

i think u understood

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

yes i did thank you

OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x − y = −3 and passes through the point (4, −2). y = −1/3x − 6 y = −1/3x − 14 y = 3x − 6 y = 3x − 14 solution: 3x-y= -3----> given writing in term of x we get y= 3x+3 here the slope = 3 now the line parallel to this line will have have the same slope thus slope = 3 point = (4, -2) equation is y+2 = 3(x-4) y= 3x-14----> right ans